OpenStudy (anonymous):

How do you get from 7 * 7^n - 3 * 3^n to 3(7^n - 3^n) + 4*7^n? This was part of the teacher's example of mathematical induction. She tried to explain this part with

OpenStudy (anonymous):

She tried to explain it with 7x - 3y = 3(x - y) + 4x. I'm just confused about the reasoning behind putting it in this form

OpenStudy (anonymous):

when writing * r u multiplying or is it an exponent?

OpenStudy (anonymous):

multiplying \[7 \times 7^n - 3 \times 3^n\]